OpenStudy (anonymous):

HELP ! Jose is paying for an eleven dollar meal using bills in his wallet. He has four one dollar bills, two five dollar bills, and two ten dollar bills. If he selects two bills at random, one at a time from his wallet, what is the probability that he will choose a ten dollar bill and a one dollar bill to pay for the meal? Show your work.

OpenStudy (johnweldon1993):

Well lets see...to start off with he has 4 + 2 + 2 = 8 bills in his wallet The probability of first choosing a 10....is 2 out of the possible 8....or 1/4 But after he takes that out, he only has 7 bills in his wallet...and the chance of taking out a 1 dollar bill....is 4/7....so the total probability will be \[\large \frac{1}{4} \times \frac{4}{7}=?\]

OpenStudy (anonymous):

1/7 ? @johnweldon1993

OpenStudy (johnweldon1993):

Correct :)